# SPHEREx: Aromatics, Aliphatics and PAH Size across the Iris Nebula

### v2

@@ 1 Introduction
The NASA Medium Class Explorer mission SpectroPhotometer for the History of the Universe, Epoch of Reionization, and Ices Explorer (SPHEREx; B. P. Crill et al., 2020) was successfully launched March 11, 2025. During its two-year mission it will be conducting the first near-infrared (IR), 0.75–5 µm, all-sky spectro-imaging survey at a spatial resolution of 6".2. Emission from polycyclic aromatic hydrocarbons (PAHs) dominates the IR emission of many astronomical objects, including interstellar medium (ISM) sources in the Milky Way, nearby galaxies, and even high-redshift galaxies (A. G. G. M. Tielens, 2008; J. Spilker et al., 2023). The blended 3.3 and 3.4 µm emission band complex is a prominent PAH feature covered by SPHEREx. Its potential to study the 3.3 and related 3.4 µm components was recently explored by E. Zhang et al. (2025) and Y.-T. Cheng et al. (2025). Obtaining 3.3 µm measurements of local interstellar medium (ISM) sources, especially those that are extended in nature, can provide valuable information on the evolution of the PAH aliphatic/aromatic ratio as well as PAH size (e.g., L. J. Allamandola et al., 1989; B. T. Draine et al., 2021). In particular with the 3.4/3.3 and 11.2/3.3 µm PAH band strength ratios, where the former traces the aliphatic/aromatic ratio and the latter is considered the ultimate tracer for PAH size (see e.g., A. Maragkoudakis et al., 2020).
@@ 2 Observations and Data Reduction
Spitzer spectral map data on NGC7023 have been taken from C. Boersma et al. (2018) and the reader is directed there for details on its reduction. The astroquery Irsa-interface was used to retrieve Multi-Extension FITS (MEF) images that have pixels overlapping the Spitzer spectral cube’s field-of-view (FOV) from the SPHEREx Archive at NASA/IPAC Infrared Science Archive (IRSA; SPHEREx Team, 2025). These data are from the SPHEREx Quick Release 2 (QR2) and consist of observations from Survey 1 with Header Data Unit (HDU) creation dates spanning 2025-09-15 to 2025-09-19.
@@ 4.2 The Maps
The maps in Figure 5 show a clear demarcation between the dense and diffuse medium towards the north and south of the FOV, respectively, that matches the morphology apparent in Figure 1. At the PDR, the emission peaks and gradually falls off when crossing into the diffuse medium and when moving towards the embedded YSO located towards the northwest.
@@
As far as I'3.4/I'3.3 is concerned, its evolution across the northwest PDR of NGC7023 is consistent with an interpretation that aliphatic hydrogen side groups have a better chance of surviving in more shielded, denser regions but, once exposed, are rapidly photo-dissociated from the parent PAH. Experiments suggest that PAHs with different side groups are released from ices at the PDR-front, where their production and residence time in the ice has promoted the addition of aliphatic side groups (C. X. Mendoza-Gomez et al., 1995; M. P. Bernstein et al., 1999, 2002)
@@
However, the right panel of Figure 7, which replaces I3.3 with '3.3, paints a different, far more chaotic picture. Gone are the clear demarcation and the smooth transition between the dense and diffuse medium when crossing the PDR-front. This difference could be due to a reduction in signal quality, but the dynamic ranges, 6.3 versus 4, remain comparable. Arguably, limitations of the spectro-photometric model used to isolate the 3.3 µm feature from the 3.4 µm sub-feature could be the cause. For example, the model described in Appendix C combines the 3.4 µm feature and underlying plateau that is shared with the 3.3 µm feature into a single Gaussian. Nevertheless, the 3.4/3.3 µm results do demonstrate merit for this approach.
@@ 4.5 The PAH Size Evolution across NGC7023
To obtain a quantitative measure for the evolution of PAH size across NGC7023, the NASA Ames PAH IR Spectroscopic Database (PAHdb hereafter; C. W. Bauschlicher et al., 2010; C. Boersma et al., 2014b; C. W. Bauschlicher et al., 2018; A. L. Mattioda et al., 2020) is employed to fit the 5.2–15 µm isolated, extinction corrected Spitzer PAH segment of the spectra. This approach is largely analogous to that used by C. Boersma et al. (2013, 2015, 2016, 2018), the main differences being: 1. the use of a much more recent version (4.00-; A. Maragkoudakis et al., 2025) of PAHdb’s library of computed PAH spectra, 2. exclusion of nitrogen containing PAHs (PANHs; per A. Ricca et al., 2024), 3. no application of an emission redshift (per C. J. Mackie et al., 2018), 4. incorporating an excitation energy of 7 eV, and 5. having Gaussian line profiles with a fixed full-width at half maximum (FWHM) of 15 cm-1. This matches the optimal PAHdb modeling configuration from A. Maragkoudakis et al. (2025). A Monte Carlo approach is taken where the spectra are varied within their uncertainties 1024 times and fitted to obtain a statistical measure, i.e., mean and standard deviation, for derived parameters. The interaction with the library, emission modeling, and fitting are done with PAHdb’s suite of Python tools222github.com/PAHdb/AmesPAHdbPythonSuite (The PAHdb Team, 2022).
Figure 8 presents maps of the determined average number of carbon atoms and small PAH fraction fsmall (N). Despite their limited span, especially for (70-76), both maps pick up the demarcation between the dense and diffuse medium towards the north and south of the FOV, respectively. As expected, increases and fsmall drops across the PDR and into the diffuse medium. Compared to B. A. Croiset et al. (2016), who find on average N in the diffuse medium and 50 at the PDR surface, the upper boundary in the present data is slightly higher (76) and does not drop below 70.
@@
The top two panels of Figure 9 present 11.2/3.3 versus . In the left and right panel the 3.3 µm PAH band strength determined through direct integration (I3.3) and spectro-photometric modeling (I'3.3) are shown, respectively. With a sigma-clip of 3 and disregarding points with a signal-to-noise ratio (SNR) of less than 3, tentative, slightly positive trend appears (R=0.54±0.05) for the former. However, use of I'3.3 reveals a less convincing negative correlation (R=-0.41±0.05). This behavior is driven by I'3.3 and reflects that seen in the right panel of Figure 7. In addition, the range in is very limited, spanning only 6 carbon atoms, despite I11.2/I'3.3 covering much of the observed astronomical variation (4-7; e.g., A. Maragkoudakis et al., 2020). In addition, careful analysis of JWST observations of the Orion Bar with PAHdb are able to recover the expected trend, albeit after averaging a large number of points (A. Maragkoudakis et al., under review). Section 4.4 provides possible explanations for the discrepancy seen here.
@@
On top of the possible reasons given in Section 4.4, it could be the intrinsic intensity of the C-H stretches and bends is a contributing factor for not observing stronger correlations. C-H stretches and bends are significantly stronger in neutral PAHs than their ionized counterparts. With larger PAHs more readily being ionized, and since the environment is highly ionized, the 11.2/3.3 µm PAH band strength ratio might not fully capture the subpopulation of large ionized PAHs.
Another possibility is that the ongoing photo-chemistry does not allow for a straightforward interpretation of PAH size, certainly not as easily as has been the case for PAH charge. Perhaps this is related to how the assorted population of freshly minted PAHs sublimated off icy grains entering the diffuse medium at the PDR-front is steadily evolving towards one of mainly fullerenes when approaching HD200775 (O. Berné & A. G. G. M. Tielens, 2012; O. Berné et al., 2013). A larger, more varied sample of sizable extended objects of combined SPHEREx and Spitzer spectral mapping observations will be needed to provide stronger evidence for the reported trends.
@@ 5 Conclusions
The 3.4/3.3 µm band strength ratio in NGC7023 shows a demarcation that matches the large-scale morphology of the region, which, in turn, is consistent with an interpretation that PAHs with aliphatic groups are able to survive in more shielded regions but get quickly removed when entering the unshielded, diffuse medium. This demarcation shows up as two distinct trends when plotting the 3.4 versus the 3.3 µm band strength, one associated with the dense medium and the other with the diffuse medium; = 0.42±0.01 and = 0.01±0.01.
@@
NASA Ames PAH IR Spectroscopic Database Monte Carlo fits to the isolated 5–15 µm Spitzer PAH spectra to retrieve the average number of carbon atoms () and small PAH fraction (; N) are able to pick up the demarcation between the dense and diffuse medium, with increasing and decreasing towards HD200775, respectively. This is despite little variation across the FOV; and .

### v3

@@ 1 Introduction
The NASA Medium Class Explorer mission SpectroPhotometer for the History of the Universe, Epoch of Reionization, and Ices Explorer (SPHEREx; B. P. Crill et al., 2020) was successfully launched March 11, 2025. During its two-year mission it will be conducting the first near-infrared (IR), 0.75–5 µm, all-sky spectro-imaging survey at a spatial resolution of 6".2. Emission from polycyclic aromatic hydrocarbons (PAHs) dominates the IR emission of many astronomical objects, including interstellar medium (ISM) sources in the Milky Way, nearby galaxies, and even high-redshift galaxies (A. G. G. M. Tielens, 2008; J. Spilker et al., 2023). The blended 3.3 and 3.4 µm emission band complex is a prominent PAH feature covered by SPHEREx. Its potential to study the 3.3 and related 3.4 µm components was recently explored by E. Zhang et al. (2025) and Y.-T. Cheng et al. (2025). Obtaining 3.3 µm measurements of local interstellar medium (ISM) sources, especially those that are extended in nature, can provide valuable information on the evolution of the PAH aliphatic/aromatic ratio as well as PAH size (e.g., L. J. Allamandola et al., 1989; A. Li & B. T. Draine, 2012; X. J. Yang et al., 2017). In particular with the 3.4/3.3 and 11.2/3.3 µm PAH band strength ratios, where the former traces the aliphatic/aromatic ratio and the latter is considered the ultimate tracer for PAH size (see e.g., A. Maragkoudakis et al., 2020).
@@ 2 Observations and Data Reduction
Spitzer spectral map data on NGC7023 have been taken from C. Boersma et al. (2018) and the reader is directed there for details on its reduction. The astroquery Irsa-interface was used to retrieve Multi-Extension FITS (MEF) images that have pixels overlapping the Spitzer spectral cube’s field-of-view (FOV) from the SPHEREx Archive at the NASA/IPAC Infrared Science Archive (IRSA; SPHEREx Team, 2025). These data are from the SPHEREx Quick Release 2 (QR2) and consist of observations from Survey 1 with Header Data Unit (HDU) creation dates spanning 2025-09-15 to 2025-09-19.
@@ 4.2 The Maps
The maps in Figure 5 show a clear demarcation between the dense and diffuse medium toward the north and south of the FOV, respectively, that matches the morphology apparent in Figure 1. At the PDR, the emission peaks and gradually falls off when crossing into the diffuse medium and when nearing the embedded YSO located toward the northwest.
@@
As far as I'3.4/I'3.3 is concerned, its evolution across the northwest PDR of NGC7023 is consistent with an interpretation that aliphatic hydrogen side groups have a better chance of surviving in more shielded, denser regions but, once exposed, are rapidly photo-dissociated from the parent PAH. This includes both superhydrogenated and methylated PAHs, where the former have been shown to produce a better match with observations (M. P. Bernstein et al., 1996; S. A. Sandford et al., 2013; X. J. Yang et al., 2020). Experiments suggest that PAHs with different side groups are released from ices at the PDR-front, where their production and residence time in the ice has promoted superhydrogenation and the addition of side groups (C. X. Mendoza-Gomez et al., 1995; M. P. Bernstein et al., 1999, 2002)
@@
However, the right panel of Figure 7, which replaces I3.3 with I'3.3, paints a different, far more chaotic picture. Gone are the clear demarcation and the smooth transition between the dense and diffuse medium when crossing the PDR-front. This difference could be due to a reduction in signal quality, but the dynamic ranges, 6.3 versus 4, remain comparable. Arguably, limitations of the spectro-photometric model used to isolate the 3.3 µm feature from the 3.4 µm sub-feature could be the cause. For example, the model described in Appendix C combines the 3.4 µm feature and underlying plateau that is shared with the 3.3 µm feature into a single Gaussian. Nevertheless, the 3.4/3.3 µm results do demonstrate merit for this approach.
@@ 4.5 The PAH Size Evolution across NGC7023
To obtain a quantitative measure for the evolution of PAH size across NGC7023, the NASA Ames PAH IR Spectroscopic Database (PAHdb hereafter; C. W. Bauschlicher et al., 2010; C. Boersma et al., 2014b; C. W. Bauschlicher et al., 2018; A. L. Mattioda et al., 2020) is employed to fit the 5.2–15 µm isolated, extinction corrected Spitzer PAH segment of the spectra. This approach is largely analogous to that used by C. Boersma et al. (2013, 2015, 2016, 2018), the main differences being: 1. the use of a much more recent version (4.00-; A. Maragkoudakis et al., 2025) of PAHdb’s library of computed PAH spectra, 2. exclusion of nitrogen containing PAHs (PANHs; per A. Ricca et al., 2024), 3. no application of an emission redshift (per C. J. Mackie et al., 2018), 4. incorporating an excitation energy of 7 eV, and 5. having Gaussian line profiles with a fixed full-width at half maximum (FWHM) of 15 cm-1. This matches the optimal PAHdb modeling configuration from A. Maragkoudakis et al. (2025). A Monte Carlo approach is taken where the spectra are varied within their uncertainties 1024 times and fitted to obtain a statistical measure, i.e., mean and standard deviation, for derived parameters222These do not capture uncertainties associated with the PAHdb spectra or the emission modeling, neither statistical nor systematic.. The interaction with the library, emission modeling, and fitting are done with PAHdb’s suite of Python tools333github.com/PAHdb/AmesPAHdbPythonSuite (The PAHdb Team, 2022).
Figure 8 presents maps of the determined average number of carbon atoms and small PAH fraction fsmall (N). Despite their limited span, especially for (70-76), both maps pick up the demarcation between the dense and diffuse medium toward the north and south of the FOV, respectively. As expected, increases and fsmall drops across the PDR and into the diffuse medium. Compared to B. A. Croiset et al. (2016), who find on average N in the diffuse medium and 50 at the PDR surface, the upper boundary in the present data is slightly higher (76) and does not drop below 70.
@@
The top two panels of Figure 9 present 11.2/3.3 versus . In the left and right panel the 3.3 µm PAH band strength determined through direct integration (I3.3) and spectro-photometric modeling (I'3.3) are shown, respectively. With a sigma-clip of 3 and disregarding points with a signal-to-noise ratio (SNR) of less than 3, a tentative, slightly positive trend appears (R=0.54±0.05) for the former. However, use of I'3.3 reveals a less convincing negative correlation (R=-0.41±0.05). This behavior is driven by I'3.3 and reflects that seen in the right panel of Figure 7. Moreover, the range in is very limited, spanning only 6 carbon atoms, despite I11.2/I'3.3 covering much of the observed astronomical variation (4-7; e.g., A. Maragkoudakis et al., 2020). In addition, careful analysis of JWST observations of the Orion Bar with PAHdb are able to recover the expected trend, albeit after averaging a large number of points (A. Maragkoudakis et al., under review). Section 4.4 provides possible explanations for the discrepancy seen here.
@@
On top of the possible reasons given in Section 4.4, it could be that the intrinsic intensity of the C-H stretches and bends is a contributing factor for not observing stronger correlations. That is, C-H stretches and bends are significantly stronger in neutral PAHs than their ionized counterparts. With larger PAHs more readily being ionized, and since the environment is highly ionized, the 11.2/3.3 µm PAH band strength ratio might not fully capture the subpopulation of large ionized PAHs.
Another possibility is that the ongoing photo-chemistry does not allow for a straightforward interpretation of PAH size, certainly not as easily as has been the case for PAH charge. Perhaps this is related to how the assorted population of freshly minted PAHs sublimated off icy grains entering the diffuse medium at the PDR-front is steadily evolving toward one of mainly fullerenes when approaching HD200775 (O. Berné & A. G. G. M. Tielens, 2012; O. Berné et al., 2013). A larger, more varied sample of sizable extended objects of combined SPHEREx and Spitzer spectral mapping observations will be needed to provide stronger evidence for the reported trends.
@@ 5 Conclusions
The 3.4/3.3 µm band strength ratio in NGC7023 shows a demarcation that matches the large-scale morphology of the region, which, in turn, is consistent with an interpretation that PAH superhydrogenation and aliphatic side groups are able to survive in more shielded regions but get quickly removed when entering the unshielded, diffuse medium. This demarcation shows up as two distinct trends when plotting the 3.4 versus the 3.3 µm band strength, one associated with the dense medium and the other with the diffuse medium; = 0.42±0.01 and = 0.01±0.01.
@@
NASA Ames PAH IR Spectroscopic Database Monte Carlo fits to the isolated 5–15 µm Spitzer PAH spectra to retrieve the average number of carbon atoms () and small PAH fraction (; N) are able to pick up the demarcation between the dense and diffuse medium, with increasing and decreasing toward HD200775, respectively. This is despite little variation across the FOV; and .